# COALA: Numerically Stable and Efficient Framework for Context-Aware Low-Rank Approximation

### v1

@@ 1 Introduction
In the context-aware approach, we consider the task of approximating a weight matrix $W\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathbb{R}^{m\times n}$ given input data $X\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathbb{R}^{n\times k}$, where $k$ is the batch size multiplied by the context length. The goal is to find a low-rank approximation $W^{\prime}$ of $W$ that maintains the performance of the neural network while reducing its computational complexity. This objective leads to the minimization of
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm. We aim to minimize $L(W^{\prime})$ with respect to the matrix $W^{\prime}\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathbb{R}^{m\times n}$, under the constraint that the rank of $W^{\prime}$ does not exceed $r$.
@@ 2 Related Work
In recent years, the compression of deep learning models has become a critical focus within the field. Several approaches have been proposed to address this challenge, including quantization [44, 22], structured pruning [27, 2, 49], and low-rank approximation methods [20, 45, 41]. Quantization allows for reducing the bit-width of model weights, thereby decreasing memory consumption and accelerating computations. Structured pruning removes unnecessary parameters and simplifies the model architecture without significant loss of accuracy. In turn, the low rank decomposition approach is often memory-efficient and can accelerate model inference, which is particularly important for tasks related to response speed and model size when deploying on mobile devices [46, 6, 25, 36, 7, 51].
@@
The work ASVD [46] proposes a solution that manages activation outliers by transforming the weight matrix based on the activation distribution. However, this solution does not achieve the best approximation error in the posed problem, providing a reasonable yet suboptimal solution [42]. Other studies, such as [42, 40][23] and [6], present solutions that attain the theoretical minimum of the error in the Frobenius norm. Nonetheless, they still rely on the formation of Gram matrices and/or inversion of small singular values.
@@ 3 Weighted Low-Rank Approximation
where $Q$ is positive definite symmetric matrix and $\operatorname{vec}\{\cdot\}$ denotes the column-wise vectorization of a matrix. By applying [28, Theorem 3] (see Appendix A) in our specific case of the matrix $Q$, we obtain:
@@
where $S=(XX^{\top})^{1/2}$ is the unique positive definite square root of $XX^{\top}$, $U\Sigma V^{\top}$ is the SVD of $WS$, and $\Sigma_{r}$ is the matrix obtained from $\Sigma$ by setting the last $n-r$ singular values to zero.
@@ 4 Inversion-Free Solution
4$[Q,R]\;\leftarrow\;\textsc{QR}(X^{\top})$triangleright Use the Tall–Skinny QR (TSQR) method, see Section 4.2.
@@ 5 Weighted Low-Rank Approximation with Regularization
So far, we have discussed various numerical aspects of solving the problem (3). However, in practice, we also want to adapt the model to fit the available examples, but not excessively, as we aim to avoid overfitting and preserve the model’s knowledge in other domains. This situation becomes particularly pronounced when data is scarce and matrix $X$ may have more columns than rows. For example, in model compression, data are often limited due to confidentiality, yet there’s a need to deploy models on devices with restricted resources [3, 26]. A similarly relevant challenge is adapting pre-trained generative models to new concepts using just a handful of images (usually 3–5) [17]. Thus, we can formulate the following minimization problem:
$\min_{\operatorname{rank}(W^{\prime})\leq k}\|WX-W^{\prime}X\|_{F}^{2}+\mu\|W^{\prime}-W\|_{F}^{2},$ (4)
@@
4Call Algorithm 1 with input $(W,X^{\prime},r)$ to compute $A$ and $B$.
5return $A$, $B$
@@
Let $W\in\mathbb{R}^{m\times n}$ and $X\in\mathbb{R}^{n\times k}$. Suppose that $X$ has $\operatorname{rank}(X)=k\geq r$ and that the singular values of $WX$ satisfy $\sigma_{r}(WX)\neq\sigma_{r+1}(WX)$, where $\sigma_{i}(\cdot)$ denotes the $i$-th largest singular value. Let $W_{0}$ denote the problem (3) $U_{r}U_{r}^{\top}W$, and let $W_{\mu}$ denote the solution to the regularized problem (4). then the following estimate holds:
$\|W_{0}-W_{\mu}\|_{F}\leq\dfrac{2\|W\|^{2}_{2}\|W\|_{F}\left(\dfrac{\sigma_{1}(X)}{\sigma_{k}(X)}+\max\left(1,\ \dfrac{\mu}{4\,\sigma_{k}^{2}(X)}\right)\right)}{\sigma_{r}^{2}\left(WX\right)-\sigma_{r+1}^{2}\left(WX\right)}\cdot\mu.$
@@ 6.1 Model compression
In this section, we evaluate the effectiveness of our regularization-based compression approach. We compare our method to approaches that do not use fine-tuning or adaptive rank selection. However, our solution can be used not only as a standalone compression technique, but also integrated into other works as part of a problem-solving framework. See the experiments in Table 234. In these tables, compression ratio refers to the memory size of the compressed model divided by the memory size of the original model. In the SVD-LLM [42] method, we replaced Cholesky factorization with a more robust version from SVD-LLM V2 [40]. We also do not post-process the obtained low-rank approximations.
Table 2: Metric values of various compression methods with the compression ratio $70\%$. Experiments were conducted using the LLaMA-3.1-8B-Instruct model on the commonsense reasoning dataset. $\text{COALA}_{\mu}$ is the regularized approach with $\mu=1$. The accuracy metrics were computed using generation.
Method PIQA WiNoG ARC-E ARC-C OBQA HSweg
Original $80.9\pm 0.9$ $73.7\pm 1.2$ $79.6\pm 0.8$ $55.0\pm 1.4$ $43.0\pm 2.0$ $79.2\pm 0.4$
ASVD $52.4{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.0}$ $48.3{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.1}$ $26.8{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.0}$ $26.1{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.3}$ $\underline{28.0}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 2.0}$ $26.0{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 0.5}$
SVD-LLM $63.1{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.0}$ $55.6{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.1}$ $45.2{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.0}$ $\underline{27.9}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.3}$ $27.4{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 2.0}$ $\underline{33.4}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 0.5}$
$\text{COALA}_{\mu=0}$ $\underline{63.2}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.0}$ $\underline{55.8}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.1}$ $\underline{45.3}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.0}$ $27.7{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.3}$ $27.6{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 2.0}$ $33.3{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 0.5}$
$\text{COALA}_{\mu}$ $\bm{66.1}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.0}$ $\bm{56.9}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.1}$ $\bm{48.6}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.0}$ $\bm{30.2}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 1.3}$ $\bm{28.8}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 2.0}$ $\bm{35.6}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm 0.5}$
@@
Table 3: Metric values of various compression methods. All computations, except for solving the weighted low-rank approximation problem, were performed in half precision (fp16). Experiments were conducted using the LLaMA-3.2-1B-Instruct model on the commonsense reasoning dataset, which was also used for validation. $\text{COALA}_{\mu}$ is regularized approach, we get $\mu=0.1$. The accuracy metrics were computed directly using the logits; thus, the process is entirely deterministic.
Ratio Method boolQ PIQA WiNoG ARC-E ARC-C OBQA HSweg
100% Original 69.5 74.4 59.5 63.2 38.1 34.6 60.7
90% ASVD 58.0 52.5 51.3 30.0 25.9 26.8 27.8
SVD-LLM 54.1 60.6 53.8 44.3 25.5 26.0 34.6
$\text{COALA}_{\mu=0}$ 57.6 60.9 53.2 43.4 27.3 26.0 34.6
$\text{COALA}_{\mu}$ 56.7 61.6 53.8 46.1 29.0 27.0 36.0
@@
Table 4: Metric values of various compression methods. Experiments were conducted using the Mistral-7B model on the WikiText2 dataset and commonsense reasoning used for validation. $\text{COALA}_{\mu}$ is regularized approach, we get $\mu=5\cdot 10^{-4}$. The results for SliceGPT [2] and FLAP [1] were taken from work [23].
@@
Ratio Method boolQ PIQA WiNoG ARC-E ARC-C OBQA HSweg
100% Original 83.9 82.0 73.9 79.5 53.9 44.0 81.0
80% ASVD 79.0 72.7 63.2 64.2 40.1 36.2 65.1
SVD-LLM 78.4 70.2 65.8 61.1 42.4 37.2 56.4
FLAP 62.2 72.3 64.0 51.0 31.9 36.8 55.9
SliceGPT 37.8 60.6 59.4 48.1 30.0 32.0 45.1
$\text{COALA}_{\mu=0}$ 78.1 70.6 65.7 61.3 38.7 36.4 56.7
$\text{COALA}_{\mu}$ 78.5 73.1 66.6 65.9 43.5 36.4 57.0
70% ASVD 74.0 65.2 58.9 49.7 33.3 32.0 45.3
SVD-LLM 74.0 65.9 60.2 52.5 36.0 32.2 45.6
FLAP 65.2 69.5 64.8 48.9 30.5 35.8 55.6
SliceGPT 37.8 54.4 51.6 35.0 22.9 26.8 32.5
$\text{COALA}_{\mu=0}$ 73.8 66.1 60.3 52.9 36.4 32.4 45.7
$\text{COALA}_{\mu}$ 74.0 66.2 60.5 53.0 39.7 32.4 45.9
@@
We conducted experiments on the models LLaMA3-8B, LLaMA3-1B [16] and Mistral-7B [5], comparing our approach with existing methods across various datasets: boolQ [8], OpenbookQA [32], WinoGrande [35], HellaSwag [47], Arc_e [9], Arc_c [10], PIQA [4]. We used A100 GPU and Tesla T4 GPU for our experiments. The results indicate that in all the considered settings our regularized algorithm systematically achieves better metrics during compression.

### v2

@@ 1 Introduction
In the context-aware approach, we consider the task of approximating a weight matrix $W\penalty 10000\ \in\penalty 10000\ \mathbb{R}^{m\times n}$ given input data $X\penalty 10000\ \in\penalty 10000\ \mathbb{R}^{n\times k}$, where $k$ is the batch size multiplied by the context length. The goal is to find a low-rank approximation $W^{\prime}$ of $W$ that maintains the performance of the neural network while reducing its computational complexity. This objective leads to the minimization of
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm. We aim to minimize $L(W^{\prime})$ with respect to the matrix $W^{\prime}\penalty 10000\ \in\penalty 10000\ \mathbb{R}^{m\times n}$, under the constraint that the rank of $W^{\prime}$ does not exceed $r$.
@@ 2 Related Work
In recent years, the compression of deep learning models has become a critical focus within the field. Several approaches have been proposed to address this challenge, including quantization [48, 24], structured pruning [30, 2, 53], and low-rank approximation methods [21, 49, 45]. Quantization allows for reducing the bit-width of model weights, thereby decreasing memory consumption and accelerating computations. Structured pruning removes unnecessary parameters and simplifies the model architecture without significant loss of accuracy. Also, the low rank decomposition approach is often memory-efficient and can accelerate model inference, which is particularly important for tasks related to response speed and model size when deploying on mobile devices [50, 6, 28, 39, 7, 55].
@@
The work ASVD [50] proposes a solution that manages activation outliers by transforming the weight matrix based on the activation distribution. However, this solution does not achieve the best approximation error in the posed problem, providing a reasonable yet suboptimal solution [46]. Other studies, such as [46, 44], [25] and [6], present solutions that attain the theoretical minimum of the error in the Frobenius norm. Nonetheless, they still rely on the formation of Gram matrices and/or inversion of small singular values.
@@ 3 Weighted Low-Rank Approximation
where $Q$ is positive definite symmetric matrix and $\operatorname{vec}\{\cdot\}$ denotes the column-wise vectorization of a matrix. By applying [31, Theorem 3] (see Appendix A) in our specific case of the matrix $Q$, we obtain:
@@
where $S=(XX^{\top})^{1/2}$ is the unique positive definite square root of $XX^{\top}$, $U\Sigma V^{\top}$ is the SVD of $WS$, and $\Sigma_{r}$ is the matrix obtained from $\Sigma$ by setting the last nrn-r singular values to zero.
@@ 4 Inversion-Free Solution
[…]
@@ 5 Weighted Low-Rank Approximation with Regularization
So far, we have discussed various numerical aspects of solving the problem (3). However, in practice, we want to adapt the model to fit the available examples, but not excessively, as we aim to avoid overfitting and preserve the model’s knowledge in other domains. This situation becomes particularly pronounced when data is scarce and matrix $X$ may have more columns than rows. For example, in model compression, data are often limited due to confidentiality, yet there’s a need to deploy models on devices with restricted resources [3, 29]. A similarly relevant challenge is adapting pre-trained generative models to new concepts using just a handful of images (usually 3–5) [17]. Thus, we can formulate the following minimization problem:
$\min_{\operatorname{rank}(W^{\prime})\leq k}\|WX-W^{\prime}X\|_{F}^{2}+\mu\|W-W^{\prime}\|_{F}^{2},$ (4)
@@
4Call Algorithm 1 with input $(W,X^{\prime},r)$ to compute AA and $B$.
5return AA, $B$
@@
Let $W\in\mathbb{R}^{m\times n}$ and $X\in\mathbb{R}^{n\times k}$. Suppose that $X$ has $\operatorname{rank}(X)=k\geq r$ and that the singular values of $WX$ satisfy $\sigma_{r}(WX)\neq\sigma_{r+1}(WX)$, where $\sigma_{i}(\cdot)$ denotes the $i$-th largest singular value. Let $W_{0}=U_{r}U_{r}^{\top}W$ denote the solution to the problem (3), and let $W_{\mu}$ denote the solution to the regularized problem (4). then the following estimate holds:
$\|W_{0}-W_{\mu}\|_{F}\leq\dfrac{2\|W\|^{2}_{2}\|W\|_{F}}{\sigma_{r}^{2}\left(WX\right)-\sigma_{r+1}^{2}\left(WX\right)}\cdot\mu.$
@@ 6.1 Model compression
In this section, we evaluate the effectiveness of our regularization-based compression approach in practice  [^1] Our code is available at https://github.com/urparkina/COALA.. We first fix the procedure for selecting the regularization parameter $\mu$. Specifically, we determine $\mu$ relative to the unregularized solution $W_{0}$ (i.e., the one obtained for $\mu=0$) according to the formula below:
$\mu=\frac{\|W_{0}X-WX\|_{F}^{2}}{\|W_{0}-W\|_{F}^{2}}\cdot\lambda,$ (5)
where $\lambda$ serves as a hyperparameter controlling the adjustment. This step is crucial because different layers of large language models exhibit substantially different norms of the weight matrices $W$, calibration matrices $X$, and their products $WX$, see, e.g., [19]. Moreover, we compare these two strategies, analyzing how the metric depends on adaptive and non-adaptive choices of $\mu$ across layers, see Figure 4. The Mistral‑7B‑Instruct model was selected due to its pronounced variation of layer‑wise norms.
Figure 4: Comparison of the impact of parameter tuning with (Equation (5)) and without considering layer‑wise norms on model quality at 70% compression, evaluated on a common‑sense reasoning dataset using the Mistral-7B‑Instruct model.
Figure 5 presents sensitivity analysis of the parameter $\lambda$, demonstrating that the optimal value of $\mu$ remains relatively stable (in the region from 1 to 10) across different settings, including various model architectures, datasets, and compression ratios.
Figure 5: The dependence of average accuracy on the parameter $\lambda$ on the commonsense reasoning dataset for different models. On the left: 70% compression ratio, on the right: 80% compression ratio.
Table 2 compares several methods without adaptive rank selection under reduced-precision (fp16) conditions. We observe that our more numerically stable formulations improve performance, with regularization providing the most consistent gains. We compare our method to approaches that do not use fine-tuning or adaptive rank selection. However, our solution can be potentially used not only as a standalone compression technique, but also integrated into other works as a part of a problem-solving framework.
Finally, Table 3 compares our approach with state-of-the-art methods that report the relevant metrics in their manuscripts. Our method achieves comparable or even superior results solely due to the use of regularization, without any additional heuristics or fine-tuning.
@@
Table 2: Metric values of various compression methods. All computations, except for solving the weighted low-rank approximation problem, were performed in half precision (fp16). Experiments were conducted using the LLaMA-3.2-1B-Instruct model compressed at 90% using text samples from the commonsense reasoning dataset, which was also used for validation.
Method boolQ PIQA WiNoG HSwag ARC-E ARC-C OBQA
Original $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}69.5_{\pm 0.7}$} $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}74.4_{\pm 1.0}$} $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}59.5_{\pm 1.3}$} $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}60.7_{\pm 0.5}$} $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}63.2_{\pm 0.9}$} $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}38.1_{\pm 2.1}$} $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}34.6_{\pm 1.4}$}
ASVD $\bm{58.0_{\pm 0.7}}$ $52.5_{\pm 1.0}$ $51.3_{\pm 1.3}$ $27.8_{\pm 0.5}$ $30.0_{\pm 0.9}$ $25.9_{\pm 2.1}$ $\underline{26.8_{\pm 1.4}}$
SVD-LLM $54.1_{\pm 0.7}$ $60.6_{\pm 1.0}$ $\underline{53.8_{\pm 1.3}}$ $34.6_{\pm 0.5}$ $\underline{44.3_{\pm 0.9}}$ $25.5_{\pm 2.1}$ $26.0_{\pm 1.4}$
$\text{COALA}_{\mu=0}$ $\underline{57.6_{\pm 0.7}}$ $\underline{60.9_{\pm 1.0}}$ $\underline{53.2_{\pm 1.3}}$ $34.6_{\pm 0.5}$ $43.4_{\pm 0.9}$ $\underline{27.3_{\pm 2.1}}$ $26.0_{\pm 1.4}$
$\text{COALA}_{\mu}$ $59.0_{\pm 0.7}$ $\bm{62.8_{\pm 1.0}}$ $\bm{54.0_{\pm 1.3}}$ $\bm{36.6_{\pm 0.5}}$ $\bm{46.2_{\pm 0.9}}$ $\bm{29.2_{\pm 2.1}}$ $\bm{27.6_{\pm 1.4}}$
@@
Table 3: Metric values of various compression methods. Experiments were conducted using the Mistral-7B model on the WikiText2 dataset and commonsense reasoning used for validation. The results for SliceGPT [2] and FLAP [1] were taken from the work [25].
@@
Ratio Method MMLU BoolQ PIQA WiNoG HSweg ARC-E ARC-C OBQA
[…]
FLAP $25.90$ 62.26{62.26} $72.31$ $64.09$ $55.94$ $51.05$ $31.91$ $36.80$
SliceGPT $28.60$ $37.86$ $60.66$ $59.43$ $45.10$ $48.15$ $30.03$ $32.00$
80% SVD-LLM $\underline{41.80}$ $\underline{68.29}$ $73.39$ $68.43$ $61.75$ $\underline{71.34}$ $\underline{40.53}$ $36.60$
SoLA $\bm{44.20}$ $66.09$ $\underline{73.67}$ $\underline{68.75}$ $\underline{63.32}$ $69.99$ $39.76$ $\underline{39.20}$
COALA $41.20$ $\bm{78.07}$ $\bm{77.04}$ $\bm{68.82}$ $\bm{65.06}$ $\bm{72.13}$ $\bm{43.43}$ $\bm{40.20}$
FLAP $26.40$ $\bm{65.26}$ $\underline{69.59}$ $\bm{64.80}$ $\bm{55.61}$ 48.91{48.91} $30.55$ 35.80{35.80}
SliceGPT $25.00$ $37.83$ $54.41$ $51.62$ $32.54$ $35.02$ $22.95$ $26.80$
70% SVD-LLM $\underline{28.20}$ $\underline{64.62}$ $64.91$ 64.17{64.17} $47.36$ $58.25$ $30.72$ $34.20$
SoLA $\bm{33.80}$ $62.57$ $68.39$ $\underline{64.48}$ $\underline{53.00}$ $\underline{60.90}$ $\underline{32.76}$ $\bm{37.60}$
COALA $27.35$ $63.82$ $\bm{70.40}$ $62.43$ 51.02{51.02} $\bm{63.63}$ $\bm{35.49}$ $\underline{36.00}$
@@
We conducted experiments on the models LLaMA3-8B, LLaMA3-1B [16] and Mistral-7B [5] (including Insrtuct versions), comparing our approach with existing methods across various datasets: boolQ [8], OpenbookQA [35], WinoGrande [38], HellaSwag [51], Arc_e [10], Arc_c [9], PIQA [4], MMLU [18]. We used A100 GPU and Tesla T4 GPU for our experiments. The results indicate that in all the considered settings our regularized algorithm systematically achieves better metrics during compression.